# Closing in on singly charged scalars

### v1

@@ Abstract
We investigate current experimental constraints and future search prospects for a hypothetical spin-zero particle that carries unit electric charge: a singly charged scalar (SCS). In addition to providing useful benchmarks for collider searches, SCS particles are also well-motivated ingredients in relatively minimal dark sectors. We focus on scenarios in which the SCS decays promptly at colliders to a lepton plus either a neutrino or an invisible dark-sector particle of negligible mass. A promptly decaying SCS can easily have appreciable branching ratios to more than one lepton flavor while remaining consistent with constraints on lepton flavor violation. This broadens the allowed range of SCS masses to extend well beyond those for right-handed selectrons, smuons, or staus. For particular benchmark SCS branching ratios, we find that SCS masses above $\sim$185 GeV and in a lower-mass window $\sim 80-125$ GeV are still allowed at 95% confidence level. We carry out Monte Carlo simulations to explore the potential of a boosted-decision-tree-based analysis to probe the surviving SCS parameter space in future searches at the (HL-)LHC, finding a significant increase in sensitivity relative to cut-based analyses both in the lower-mass window and at higher SCS masses.
@@ I Introduction
The Large Hadron Collider (LHC) has proven to be a remarkably powerful probe of physics beyond the Standard Model (SM). So far, this has meant strong constraints on beyond-the-SM (BSM) scenarios, particularly those with missing-energy signatures. Using supersymmetry (SUSY) as a benchmark, masses beyond 2 TeV have been probed for gluinos Aad et al. (2021a); Sirunyan et al. (2019a) and up to about 1 TeV for electroweak charged superpartners Aad et al. (2021b); Tumasyan et al. (2023a). In this paper, we focus on the phenomenology of a hypothetical singly charged scalar (SCS), a BSM particle whose properties are much less constrained Cao et al. (2018); Crivellin et al. (2021). We aim to motivate collider searches dedicated to probing the SCS parameter space to the fullest extent possible, including currently untested regions that extend below 100 GeV in SCS mass.
@@
[…]
  • For ${\mathcal{B}}_{\mu}=100\%$, the LEP exclusion is $M_{\Phi}<97.3$ GeV, which overlaps with the combined LHC exclusion, $45093.5<M_{\Phi}<$ 450 GeV Aad et al. (2014, 2020), leaving no gap at 95% CL.
@@ IV Reinterpretation of an ATLAS slepton search
Figure 2: Reinterpretation of Ref. Aad et al. (2020), the ATLAS search for direct slepton production, for an SCS with ${\mathcal{B}}_{e}={\mathcal{B}}_{\mu}=25\%$ and ${\mathcal{B}}_{\tau}=50\%$. The contours give $\mu_{95}$, the 95% CL upper limit on the signal strength, as a function of SCS mass; masses with $\mu_{95}<1$ are ruled out at 95% CL for the chosen branching ratios. The blue and red contours apply, respectively, for an SCS that decays through the DM coupling of Eq. (2), and for an SCS decays through the neutrino coupling of Eq. (1). The two scenarios give slightly different results because the helicity of $\tau$s produced in SCS decays impacts the $e/\mu$ spectrum from leptonic $\tau$ decays. Not including events with SCS decays to $\tau$ gives the dashed, green contour.
@@
We present signal-strength upper limits for an SCS with our benchmark branching ratios in Figure 2, which shows that only the mass range $125\text{ GeV}\lesssim M_{\Phi}\lesssim 185\text{ GeV}$ is excluded at 95% CL. The relatively weak limits on an SCS with appreciable ratios to more than one lepton flavor motivate future dedicated collider searches for this scenario. Our results are roughly consistent with the lower-bound estimate $M_{\Phi}\gtrsim 200$ GeV obtained by the authors of Ref. Crivellin et al. (2021), based on cross-section times branching ratio alone (i.e. without simulating signal efficiencies) and considering only same-flavor events. Their estimate does not fully capture the $M_{\Phi}$ dependence of the signal efficiency, which opens a window at lower masses, $M_{\Phi}\lesssim 125$ GeV. We present results for other simple sets of branching ratios in Appendix C. For a two-flavor scenario with ${\mathcal{B}}_{\mu}={\mathcal{B}}_{\tau}=50\%$, for example, we estimate that the mass range excluded by the ATLAS search expands to $120\text{ GeV}\lesssim M_{\Phi}\lesssim 255\text{ GeV}$.
@@ V.1 Other LHC searches for selectrons and smuons
A CMS search Sirunyan et al. (2021) based on 137 fb-1 places constraints on direct slepton pair production that are comparable to those obtained by the ATLAS search used in Sec. IV, Ref. Aad et al. (2020). For example, both searches place an upper limit of $\sim 2$ fb on the total production cross section for 200-GeV degenerate sleptons (${\tilde{e}}_{L,R}$ and ${\tilde{\mu}}_{L,R}$) decaying to leptons plus a massless neutralino. The CMS analysis defines signal regions to allow SF events only, while DF events are steered into control regions. This would appear to reduce the sensitivity of the CMS search to our SCS benchmark. More recently, a CMS Physics Analysis Summary CMS presents preliminary results on a search for new physics in events with oppositely charged leptons and missing transverse momentum, using both SF and DF control regions. We have not attempted to carry out a detailed reinterpretation of this study, but we have checked that if we estimate the SCS to have a signal efficiency smaller than that that for degenerate sleptons by a factor of four (to approximately take into account, for our SCS benchmark, the requirement of two light leptons, whether SF or DF), the signal-strength upper limits come out to 2.7 for $M_{\Phi}=120$ GeV and 2.3 for $M_{\Phi}=180$ GeV, suggesting that this analysis is not likely to shrink the allowed mass ranges identified in Fig. 2. For these reasons, and because the supplementary materials provided with Ref. Aad et al. (2020) make that analysis particularly straightforward to reinterpret for an SCS that can decay to multiple flavors, we have chosen to focus on the ATLAS search.
In the lower-mass region, an ATLAS 8 TeV search Aad et al. (2014) places stronger constraints on right-handed sleptons than the 13 TeV search ATLAS search, exluding ${\tilde{e}}_{R}$ and ${\tilde{\mu}}_{R}$ for masses down to 107.4 GeV and 93.5 GeV, respectively, for the case of a massless neutralino hd (6). To obtain a rough estimate of signal-strength upper limits for our SCS benchmark, we again approximate the cross section upper limit for the SCS to be the reported cross section upper limit for combined ${\tilde{e}}_{R}$ and ${\tilde{\mu}}_{R}$ production divided by four, because only about a quarter of the signal events have two light leptons (and DF events are included in the signal regions for this analysis). Focusing on the lower-mass region, this leads to signal-strength upper limits of $\simeq 2.4$ and 4.5 for SCS masses of 125 GeV and 100 GeV, respectively. Comparing with Fig. 2, the constraint we obtained from the 13 TeV search is stronger at 125 GeV. The 8 TeV constraint is stronger at 100 GeV, but does not qualitatively change the basic point that there exists surviving parameter space to be probed.
@@ V.4 Low-energy probes and precison electroweak data
As discussed, for example, in Ref. Cao et al. (2018), precision electroweak constraints on an SCS are relatively weak. Applying the formulas of Ref. Grimus et al. (2008) to an SCS, one finds that the oblique parameters $S$ and $U$ are never larger in magnitude than $\sim 5\times 10^{-3}$ even for SCS masses going down to $M_{Z}/2$, while $T$ does not receive a one-loop correction. Because corrections to oblique parameters of order a few ×$\times 10^{-2}$ are consistent with existing data, we conclude that the SCS parameter space is not significantly constrained by precision electroweak data for SCS masses above $M_{Z}/2$.
@@ VI Potential of future BDT-based searches
Unlike the recast of Sec. IV, this study requires our own MC simulations of SM background processes. In the fit to data performed in Ref. Aad et al. (2020), the four backgrounds $WW$, $WZ$, $ZZ$, and $t{\overline{t}}$ account for 90% of the fitted background in the combined SF0J/SF1J/DF0J/DF1J signal regions defined by ATLAS (which have $M_{T_{2}}>100$ GeV), with the following breakdown: 56% for $WW$, 6% for $WZ$, 14% for $ZZ$, and 14% for $t{\overline{t}}$. We use the same MC pipeline described for the SCS signal in Sec. IV to simulate these four main backgrounds; we relegate further details about our background MC simulations to Appendix. A.
@@ VII Conclusions
We have investigated the collider phenomenology of a singly charged scalar that decays promptly to lepton plus invisible. We have seen the challenge of probing the full SCS parameter space: SCS masses above $\sim 185$ GeV still allowed for the benchmark SCS branching ratios we adopted for our study, along with a lower-mass window around $\sim 100$ GeV. Appropriately designed dedicated searches leveraging higher integrated luminosity will be able to make significant inroads addressing the surviving parameter space, and they seem well motivated given the relative simplicity of this BSM extension. Experimental searches should of course aim to probe the SCS parameter space as broadly as possible, including various branching ratio scenarios, and a range of masses for the invisible particle, which we have set to zero here for simplicity. Moving forward, searches in the $\tau$ and $e/\mu$ channels will continue to be complementary probes of the SCS parameter space. In the event of a discovery, it will be necessary to search in as many flavor combinations as possible to measure ${\mathcal{B}}_{e,\mu,\tau}$. We leave for future work the question of whether $e\tau$ and $\mu\tau$ events might be useful as SCS probes despite the large background from $W$+jets with a fake hadronic $\tau$. Looking beyond the LHC, future lepton colliders would provide powerful probes of SCS extensions of the SM.
@@ Appendix A Monte Carlo event generation
In this Appendix we provide details on our MC simulations of the SCS signal and of the $WW$, $WZ$, $ZZ$, and $t{\overline{t}}$ SM background process. The SCS reintepretaion of Ref. Aad et al. (2020) (Sec. IV) uses only our SCS MC samples, while the analyses of Sec. VI require our background MC samples in addition. For all simulated processes, we use MadGraph/MadEvent Alwall et al. (2014) for parton-level event generation, Pythia8 Bierlich et al. (2022) for hadronization and showering, and Delphes de Favereau et al. (2014) to simulate detector effects. We adopt MLM jet matching Mangano et al. (2007) with up to two additional jets at the matrix element level. We set MadEvent run card parameter xqcut, which determines the minimum $k_{T}$ between partons at the matrix-element level, to be 1/3 the mass of the SCS particle for signal simulations, at 30 GeV for $WW$, $WZ$, and $ZZ$ production, and at 60 GeV for $t{\overline{t}}$ production. In each case we take the MLM matching scale $Q_{\text{c}ut}$ to be 3/2 times xqcut.
@@
For the BDT-based analysis of Sec. VI, we instead require only $M_{ll}>100$ GeV and $M_{T_{2}}>85$ GeV, and we bin in the BDT variable, with twenty equal-width bins covering the full range from $-1$ to $+1$.
@@
In our background MC simulations, we force all $W$ and $Z$ bosons to decay leptonically, with all three lepton flavors included (this includes both $W$’s in $t{\overline{t}}$ events). To generate the $ZZ$ sample, one $Z$ is forced to decay to neutrinos and the other is forced to decay to charged leptons, while for the $WZ$ sample the $Z$ is decayed to charged leptons. By making these choices we effectively require at least two leptons and missing transverse momentum at parton level. For all four samples we allow one (but not both) of the $W/Z/t$ particles to be produced off-shell, where the kinematic separation between on- and off-shell is set by the value of the relevant MadEvent parameter bwcutoff $15$, meaning the invariant mass of the $W/Z/t$ decay products is restricted to deviate from the pole mass by no more than ±$\pm 15\times\Gamma_{W/Z/t}$. By off-shell $Z$, we simply mean a $l^{+}l^{-}$ or $\nu{\overline{\nu}}$ pair with invariant mass far from $M_{Z}$; in the charged lepton case, both $\gamma$ and $Z$ contributions are included at the amplitude level, with the dilepton invariant mass restricted to $m_{ll}>60$ GeV. For the SCS signal, we assume that the SCS width is small enough to neglect off-shell production and interference between SCS and SM processes. Our small-width assumption is motivated in part by constraints on lepton flavor violation, although, as discussed in Sec. V.4, those constraints depend on the flavor structure of the SCS couplings.
@@ Appendix B Recast cross checks
As a check on the methods we employ in Section IV for our SCS reinterpretation of Ref. Aad et al. (2020), we repeat the exercise for a simplified SUSY model with degenerate sleptons (${\tilde{e}}_{L,R}$ and ${\tilde{\mu}}_{L,R}$) and an effectively massless lightest neutralino. The supplementary HistFactory materials provided by ATLAS include expected counts in all signal regions for a variety of slepton masses for this model. We generate our own MC samples for direct slepton production using the same settings described in Appendix A for SCS pair production. We compare the signal-strength upper limits we obtain using our signal MC with the limits that ATLAS’s MC gives in Fig. 4. The HistFactory files provided by ATLAS include systematic and MC-statistical uncertainties for their slepton signals. To disentangle whether discrepancies are due to these signal uncertainties (which we do not attempt to incorporate for our MC signal), or due to differences in our simulations, we make the comparison with and without the inclusion of the signal uncertainties reported by ATLAS. Note that the 700 GeV slepton mass lower bound implied by the blue contour in Fig. 4(a) is consistent with the lower bound reported in Ref. Aad et al. (2020). In the $200-500$ GeV mass range, we see that our signal MC gives agreement with ATLAS results at a better than $10\%$ level. We get a somewhat larger discrepancy, approaching $20\%$, for a slepton mass of 100 GeV. However, it seems that the ATLAS analysis is nowhere near sensitive to our benchmark SCS model at that mass, as Fig. 2 shows. Unfortunately, ATLAS does not provide information for simplified SUSY models with a near-massless neutralino and slepton masses between 100 and 200 GeV.

### v2

@@ Abstract
We investigate current experimental constraints and future search prospects for a hypothetical $SU(2)_{w}$ and $SU(3)_{c}$ singlet spin-zero particle that carries unit electric charge: a singly charged scalar (SCS). In addition to providing useful benchmarks for collider searches, SCS particles are also well-motivated ingredients in relatively minimal dark sectors. We focus on scenarios in which the SCS decays promptly at colliders to a lepton plus either a neutrino or an invisible dark-sector particle of negligible mass. A promptly decaying SCS can easily have appreciable branching ratios to more than one lepton flavor while remaining consistent with constraints on lepton flavor violation. This broadens the allowed range of SCS masses to extend well beyond those for right-handed selectrons, smuons, or staus. For particular benchmark SCS branching ratios, we find that SCS masses above $\sim$185 GeV and in a lower-mass window $\sim 80-125$ GeV are still allowed at 95% confidence level. We carry out Monte Carlo simulations to explore the potential of a boosted-decision-tree-based analysis to probe the surviving SCS parameter space in future searches at the (HL-)LHC, finding a significant increase in sensitivity relative to cut-based analyses both in the lower-mass window and at higher SCS masses.
@@ I Introduction
The Large Hadron Collider (LHC) has proven to be a remarkably powerful probe of physics beyond the Standard Model (SM). So far, this has meant strong constraints on beyond-the-SM (BSM) scenarios, particularly those with missing-energy signatures. Using supersymmetry (SUSY) as a benchmark, masses beyond 2 TeV have been probed for gluinos Aad et al. (2021a); Sirunyan et al. (2019a) and up to about 1 TeV for electroweak charged superpartners Aad et al. (2021b); Tumasyan et al. (2023a). In this paper, we focus on the phenomenology of a hypothetical singly charged scalar (SCS), a BSM particle whose properties are much less constrained Cao et al. (2018); Babu et al. (2020); Crivellin et al. (2021). We aim to motivate collider searches dedicated to probing the SCS parameter space to the fullest extent possible, including currently untested regions that extend below 100 GeV in SCS mass.
@@
  • For ${\mathcal{B}}_{e}=100\%$, LHC searches exclude the mass range $107$ $\lesssim M_{\Phi}\lesssim 425 GeV$ Aad et al. (2014, 2020), leaving a small gap from $\sim 100-107$ GeV between LEP and LHC exclusions. Our estimate of the lower boundary of this mass window is only approximate, because the $e^{+}e^{-}\rightarrow{\tilde{e}}_{R}{\tilde{e}}_{R}^{*}$ amplitude includes a contribution from $t$-channel neutralino exchange (calculated in Ref. lep for a particular SUSY model), which is absent in the SCS scenario with sufficiently small $F$ couplings.
  • For ${\mathcal{B}}_{\mu}=100\%$, the LEP exclusion is $M_{\Phi}<97.3$ GeV, which overlaps with the combined LHC exclusion, $93.5<M_{\Phi}<450$ GeV Aad et al. (2014, 2020), leaving no gap at 95% CL.
@@ IV Reinterpretation of an ATLAS slepton search
[…]
We present signal-strength upper limits for an SCS with our benchmark branching ratios in Figure 2, which shows that only the mass range $125\text{ GeV}\lesssim M_{\Phi}\lesssim 185\text{ GeV}$ is excluded at 95% CL. The relatively weak limits on an SCS with appreciable ratios to more than one lepton flavor motivate future dedicated collider searches for this scenario.
We obtain stronger limits than the LHC constraints from Ref. Cao et al. (2018), which used Run I results, and from Ref. Babu et al. (2020), which used a Run II CMS search based on an integrated luminosity of 36 fb-1 Sirunyan et al. (2019b). Our results are roughly consistent with the lower-bound estimate $M_{\Phi}\gtrsim 200$ GeV obtained in Ref. Crivellin et al. (2021) using the same ATLAS search we consider here. The authors of Ref. Crivellin et al. (2021) arrived at their estimate by using cross-section times branching ratio alone (i.e. without simulating signal efficiencies) and considering only same-flavor events. Their estimate does not fully capture the $M_{\Phi}$ dependence of the signal efficiency, which opens a window at lower masses, $M_{\Phi}\lesssim 125$ GeV.
We present results for other simple sets of branching ratios in Appendix C. For a two-flavor scenario with ${\mathcal{B}}_{\mu}={\mathcal{B}}_{\tau}=50\%$, for example, we estimate that the mass range excluded by the ATLAS search expands to $120\text{ GeV}\lesssim M_{\Phi}\lesssim 255\text{ GeV}$.
@@ V.1 Other LHC searches for selectrons and smuons
A CMS search Sirunyan et al. (2021) based on 137 fb-1 places constraints on direct slepton pair production that are comparable to those obtained by the ATLAS search used in Sec. IV, Ref. Aad et al. (2020). For example, both searches place an upper limit of $\sim 2$ fb on the total production cross section for 200-GeV degenerate sleptons (${\tilde{e}}_{L,R}$ and ${\tilde{\mu}}_{L,R}$) decaying to leptons plus a massless neutralino. The CMS analysis defines signal regions to allow SF events only, while DF events are steered into control regions. This would appear to reduce the sensitivity of the CMS search to our SCS benchmark. More recently, a CMS Physics Analysis Summary CMS presents preliminary results on a search for new physics in events with oppositely charged leptons and missing transverse momentum, using both SF and DF control regions. We have not attempted to carry out a detailed reinterpretation of this study, but we have checked that if we estimate the SCS to have a signal efficiency smaller than that for degenerate sleptons by a factor of four (to approximately take into account, for our SCS benchmark, the requirement of two light leptons, whether SF or DF), the signal-strength upper limits come out to 2.7 for $M_{\Phi}=120$ GeV and 2.3 for $M_{\Phi}=180$ GeV, suggesting that this analysis is not likely to shrink the allowed mass ranges identified in Fig. 2. For these reasons, and because the supplementary materials provided with Ref. Aad et al. (2020) make that analysis particularly straightforward to reinterpret for an SCS that can decay to multiple flavors, we have chosen to focus on the ATLAS search.
In the lower-mass region, an ATLAS 8 TeV search Aad et al. (2014) places stronger constraints on right-handed sleptons than the 13 TeV search ATLAS search, excluding ${\tilde{e}}_{R}$ and ${\tilde{\mu}}_{R}$ for masses down to 107.4 GeV and 93.5 GeV, respectively, for the case of a massless neutralino hd (6). To obtain a rough estimate of signal-strength upper limits for our SCS benchmark, we again approximate the cross section upper limit for the SCS to be the reported cross section upper limit for combined ${\tilde{e}}_{R}$ and ${\tilde{\mu}}_{R}$ production divided by four, because only about a quarter of the signal events have two light leptons (and DF events are included in the signal regions for this analysis). Focusing on the lower-mass region, this leads to signal-strength upper limits of $\simeq 2.4$ and 4.5 for SCS masses of 125 GeV and 100 GeV, respectively. Comparing with Fig. 2, the constraint we obtained from the 13 TeV search is stronger at 125 GeV. The 8 TeV constraint is stronger at 100 GeV, but does not qualitatively change the basic point that there exists surviving parameter space to be probed. Similarly, an ATLAS Run-II slepton search based on 36.1 fb-1, Ref. Aaboud et al. (2018), does not further impact the parameter space for our SCS benchmark. Using the auxiliary material from Ref. 81996hd (819) to perform the identical exercise as for the ATLAS 8 TeV search, we estimate signal strength upper limits of $\sim 2$ and $\sim 30$ for SCS masses of 150 GeV and 100 GeV, respectively.
@@ V.4 Low-energy probes and precison electroweak data
As discussed, for example, in Ref. Cao et al. (2018), precision electroweak constraints on an SCS are relatively weak. Applying the formulas of Ref. Grimus et al. (2008) to an SCS, one finds that the oblique parameters $S$ and $U$ are never larger in magnitude than $\sim 5\times 10^{-3}$ even for SCS masses going down to $M_{Z}/2$, while $T$ does not receive a one-loop correction. Because corrections to oblique parameters of order a few $\times 10^{-2}$ are consistent with existing data, we conclude that the SCS parameter space is not significantly constrained by precision electroweak data for SCS masses above $M_{Z}/2$.
@@ VI Potential of future BDT-based searches
Unlike the recast of Sec. IV, this study requires our own MC simulations of SM background processes. In the fit to data performed in Ref. Aad et al. (2020), the four backgrounds $WW$, WZWZ, ZZZZ, and $t{\overline{t}}$ account for 90% of the fitted background in the combined SF0J/SF1J/DF0J/DF1J signal regions defined by ATLAS (which have $M_{T_{2}}>100$ GeV), with the following breakdown: 56% for $WW$, 6% for $WZ$, 14% for $ZZ$, and 14% for $t{\overline{t}}$. We use the same MC pipeline described for the SCS signal in Sec. IV to simulate these four main backgrounds; we relegate further details about our background MC simulations to Appendix. A.
@@ VII Conclusions
We have investigated the collider phenomenology of a singly charged scalar that decays promptly to lepton plus invisible. We have seen the challenge of probing the full SCS parameter space: SCS masses above $\sim 185$ GeV still allowed for the benchmark SCS branching ratios we adopted for our study, along with a lower-mass window around $\sim 100$ GeV. Appropriately designed dedicated searches leveraging higher integrated luminosity will be able to make significant inroads addressing the surviving parameter space, and they seem well motivated given the relative simplicity of this BSM extension. Experimental searches should of course aim to probe the SCS parameter space as broadly as possible, including various branching ratio scenarios, and a range of masses for the invisible particle, which we have set to zero here for simplicity. Moving forward, searches in the $\tau$ and $e/\mu$ channels will continue to be complementary probes of the SCS parameter space. In the event of a discovery, it will be necessary to search in as many flavor combinations as possible to measure ${\mathcal{B}}_{e,\mu,\tau}$. We leave for future work the question of whether $e\tau$ and $\mu\tau$ events might be useful as SCS probes despite the large background from WW+jets with a fake hadronic $\tau$. Looking beyond the LHC, future lepton colliders would provide powerful probes of SCS extensions of the SM.
@@ Appendix A Monte Carlo event generation
In this Appendix we provide details on our MC simulations of the SCS signal and of the $WW$, WZWZ, ZZZZ, and $t{\overline{t}}$ SM background process. The SCS reintepretaion of Ref. Aad et al. (2020) (Sec. IV) uses only our SCS MC samples, while the analyses of Sec. VI require our background MC samples in addition. For all simulated processes, we use MadGraph/MadEvent Alwall et al. (2014) for parton-level event generation, Pythia8 Bierlich et al. (2022) for hadronization and showering, and Delphes de Favereau et al. (2014) to simulate detector effects. We adopt MLM jet matching Mangano et al. (2007) with up to two additional jets at the matrix element level. We set MadEvent run card parameter xqcut, which determines the minimum $k_{T}$ between partons at the matrix-element level, to be 1/3 the mass of the SCS particle for signal simulations, at 30 GeV for $WW$, WZWZ, and $ZZ$ production, and at 60 GeV for $t{\overline{t}}$ production. In each case we take the MLM matching scale $Q_{\text{c}ut}$ to be 3/2 times xqcut.
@@
For the BDT-based analysis of Sec. VI, we instead require only $M_{ll}>100$ GeV and $M_{T_{2}}>85$ GeV, and we bin in the BDT variable, with twenty equal-width bins covering the full range from 1-1 to +1+1.
@@
In our background MC simulations, we force all $W$ and $Z$ bosons to decay leptonically, with all three lepton flavors included (this includes both $W$’s in $t{\overline{t}}$ events). To generate the $ZZ$ sample, one $Z$ is forced to decay to neutrinos and the other is forced to decay to charged leptons, while for the $WZ$ sample the $Z$ is decayed to charged leptons. By making these choices we effectively require at least two leptons and missing transverse momentum at parton level. For all four samples we allow one (but not both) of the $W/Z/t$ particles to be produced off-shell, where the kinematic separation between on- and off-shell is set by the value of the relevant MadEvent parameter bwcutoff $15$, meaning the invariant mass of the $W/Z/t$ decay products is restricted to deviate from the pole mass by no more than $\pm 15\times\Gamma_{W/Z/t}$. By off-shell $Z$, we simply mean a $l^{+}l^{-}$ or $\nu{\overline{\nu}}$ pair with invariant mass far from $M_{Z}$; in the charged lepton case, both $\gamma$ and $Z$ contributions are included at the amplitude level, with the dilepton invariant mass restricted to $m_{ll}>60$ GeV. For the SCS signal, we assume that the SCS width is small enough to neglect off-shell production and interference between SCS and SM processes. Our small-width assumption is motivated in part by constraints on lepton flavor violation, although, as discussed in Sec. V.4, those constraints depend on the flavor structure of the SCS couplings.
@@ Appendix B Recast cross checks
As a check on the methods we employ in Section IV for our SCS reinterpretation of Ref. Aad et al. (2020), we repeat the exercise for a simplified SUSY model with degenerate sleptons (${\tilde{e}}_{L,R}$ and ${\tilde{\mu}}_{L,R}$) and an effectively massless lightest neutralino. The supplementary HistFactory materials provided by ATLAS include expected counts in all signal regions for a variety of slepton masses for this model. We generate our own MC samples for direct slepton production using the same settings described in Appendix A for SCS pair production. We compare the signal-strength upper limits we obtain using our signal MC with the limits that ATLAS’s MC gives in Fig. 4. The HistFactory files provided by ATLAS include systematic and MC-statistical uncertainties for their slepton signals. To disentangle whether discrepancies are due to these signal uncertainties (which we do not attempt to incorporate for our MC signal), or due to differences in our simulations, we make the comparison with and without the inclusion of the signal uncertainties reported by ATLAS. Note that the 700 GeV slepton mass lower bound implied by the blue contour in Fig. 4(a) is consistent with the lower bound reported in Ref. Aad et al. (2020). In the 200500200-500 GeV mass range, we see that our signal MC gives agreement with ATLAS results at a better than $10\%$ level. We get a somewhat larger discrepancy, approaching $20\%$, for a slepton mass of 100 GeV. However, it seems that the ATLAS analysis is nowhere near sensitive to our benchmark SCS model at that mass, as Fig. 2 shows. Unfortunately, ATLAS does not provide information for simplified SUSY models with a near-massless neutralino and slepton masses between 100 and 200 GeV.